OpenStudy (anonymous):

name five fractions that are near but not equivalent to 4/12

zepdrix (zepdrix):

near is pretty subjective :) But let's start with the easy ones. \[\Large \frac{5}{12},\quad \frac{3}{12}\]Those are relatively close to 4/12, yes?

OpenStudy (anonymous):

like that it not equivalent to 4/12 but near to it

OpenStudy (anonymous):

zepredix ok that is tat is the first step so what do next

zepdrix (zepdrix):

Do you understand how \(\Large \dfrac{8}{24}\) is equivalent to \(\Large \dfrac{4}{12}\) ? I multiplied the top and bottom by 2, so it's still `equivalent`. We want to use a big denominator like this, so we can take smaller pieces out of our 4/12.

OpenStudy (anonymous):

i get that but wht fractions r near but not equivalent to 4/12 and i need 5 please help me

zepdrix (zepdrix):

Like maybe you'll notice that 3/12 is off from 4/12 by 1/12. That's kind of a lot. If instead we took 1 away from our 8/24, giving us 7/24, we're now only 1/24 away from 4/12, a bit closer!

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

So we can't use 8/24 because it is equivalent. But just add and subtract 1 from the numerator, that should give you a couple numbers close to 4/12

OpenStudy (anonymous):

continue the lesson

zepdrix (zepdrix):

So we need a 5th? Hmm

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are u a teacher

OpenStudy (anonymous):

and if u r not what grade r u in

zepdrix (zepdrix):

We can take our 4/12 and multiply the top and bottom by 4,\[\Large \frac{16}{48}\]This again is equivalent to 4/12, so we can't use it. We'll have to use numbers close to it. A teacher? :) hah, no. I do some tutoring though.

zepdrix (zepdrix):

I'm in my 3rd year of college :o

OpenStudy (anonymous):

cool so i get that part and what do u want to be later on in life and you r smart and what college do u go to i

zepdrix (zepdrix):

I'm not sure yet :) still trying to figure that out. I want to make stuff. Lots of stuff. Invent stuff: Design computer hardware and peripherals, cars, airplanes, furniture.. I dunno stuff. Gotta spend some more time in school before I know for sure though.

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

i want to be a architectural engineering

zepdrix (zepdrix):

ooo neato :D

OpenStudy (anonymous):

lol so lets get back to the problem

OpenStudy (anonymous):

so where were we at

zepdrix (zepdrix):

\[\Large \frac{8}{24} \text{ and }\frac{16}{48}\text{ are equivalent to }\frac{4}{12}\]

zepdrix (zepdrix):

So we want to use these new fractions, and add or subtract 1 from the numerator to get something very close to our equivalence.

OpenStudy (anonymous):

yeh so we cant use it

OpenStudy (anonymous):

ok continue

zepdrix (zepdrix):

So what is a number close to 16/48?

OpenStudy (anonymous):

ummmmm i 1/3 because i simplify the number is it correct

zepdrix (zepdrix):

No we don't want to simplify the number silly :o That will just give us another number that is equivalent to 4/12. We want to adjust the numerator. So how about 17/48? That's pretty close to 16/48, right?

OpenStudy (anonymous):

oooooooooooooooooooooookkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkk i get it and 17/48 is pretty close to 16/48

zepdrix (zepdrix):

and 15/48 is also close.

zepdrix (zepdrix):

So I think that gives us 6 numbers to work with.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

Oh also, \(\Large \color{royalblue}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so where were we at. so what r the numbers that are near but not equivalent to 4/12

OpenStudy (anonymous):

hello

OpenStudy (mathstudent55):

\( \dfrac{4}{12} = \dfrac{8}{24} = \dfrac{12}{36} = \dfrac{16}{48} = \dfrac{20}{60} = \dfrac{24}{72}\) \(\dfrac{4}{12} \approx \dfrac{8}{25} \) \(\dfrac{4}{12} \approx \dfrac{12}{37} \) \(\dfrac{4}{12} \approx \dfrac{16}{49} \) \(\dfrac{4}{12} \approx \dfrac{20}{61} \) \(\dfrac{4}{12} \approx \dfrac{24}{73} \)